# Lorentzian path-integral of Robin Universe

### v1

@@ II Mini-superspace action and Boundary Terms
Interestingly, for Weyl-flat metrics there is an identity relating the Riemann tensor with the Ricci-tensor and the Ricci scalar. This is given by,
@@
This is very useful relation as when it is utilized in the GB-gravity action then we get the following for case of Weyl-flat metrics
@@
where $V_{D-1}$ is the volume of $D-1$ dimensional sphere and is given by,
@@
Lapse and scale-factor rescaling allows one to recast this action into a more simpler form. This change of variables is given by
@@
Under this transformation the gravitational action becomes the following
@@
In this the $(\dot{})$ represent time $t$-derivative. For some terms one can do the integration by parts which allow us to simplify a bit more and combine some of the terms. This eventually leads to
@@
Notice that there are two total derivative terms: one arises from the EH-part of gravitational action, while the other is the GB term itself which is topological in $4D$. Note also that in the following, we will make use of the convention $V_{3}=8\pi G_{N}$.
At this point one can setup the variational problem by varying the action given in eq. (19) with respect to $q(t)$, which we do it by working in the ADM gauge $\dot{N}=0$ (setting $N(t)=N_{c}$, a constant). This can be implemented systematically by writing
@@
where one assumes $\bar{q}(t)$ to satisfy the equation of motion, $\delta q(t)$ is the fluctuation around this. The parameter $\epsilon$ is inserted to keep track of terms in expansion, which becomes handy as one goes to higher-orders in perturbative expansion.
@@
Note that in the above there are two total time-derivative terms. These eventually become boundary terms. For a consistent variational problem, each of these must either vanish for the given choice of boundary conditions or need to be appropriately cancelled by the addition of extra boundary terms. On the other hand, the collection of terms proportional to $\delta q$ gives the equation of motion for $q$,
@@
$S_{\rm bdy}=\frac{\epsilon}{2}\biggl{[}\frac{3}{N_{c}}\left(\bar{q}_{f}\delta\dot{q}_{f}-\bar{q}_{i}\delta\dot{q}_{i}\right)+24\alpha\left\{\left(\frac{k\delta\dot{q}_{f}}{2N_{c}}+\frac{\dot{\bar{q}}_{f}^{2}\delta\dot{q}_{f}}{8N_{c}^{3}}\right)-\left(\frac{k\delta\dot{q}_{i}}{2N_{c}}+\frac{\dot{\bar{q}}_{i}^{2}\delta\dot{q}_{i}}{8N_{c}^{3}}\right)\right\}\biggr{]}\,,$ (25)
@@
$S_{\rm bdy}=-\epsilon\biggl{[}\left(\bar{q}_{f}\delta\pi_{f}-\bar{q}_{i}\delta\pi_{i}\right)+4\alpha\left\{\left(k\delta\pi_{f}+\frac{\bar{\pi}_{f}^{2}\delta\pi_{f}}{27}\right)-\left(k\delta\pi_{i}+\frac{\bar{\pi}_{i}^{2}\delta\pi_{i}}{27}\right)\right\}\biggr{]}\,.$ (28)
@@ II.2 Robin Boundary condition (RBC) at $t=0$
.$\biggl{.}S_{\rm bdy}\biggr{|}_{\rm RBC}=-\epsilon\biggl{[}\bar{q}_{f}\delta\pi_{f}+\beta\bar{q}_{i}\delta q_{i}+4\alpha\left\{k\delta\pi_{f}+\frac{\bar{\pi}_{f}^{2}\delta\pi_{f}}{9}+\beta\left(k\delta q_{i}+\frac{(\bar{P}_{i}-\beta\bar{q}_{i})^{2}\delta q_{i}}{9}\right)\right\}\biggr{]}\,.$ (36)
@@ III Transition Amplitude
$G[{\rm Bd}_{f},{\rm Bd}_{i}]=\int_{0^{+}}^{\infty}{\rm d}N_{c}\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}{\cal D}q(t)\,\,\exp\left(\frac{i}{\hbar}S_{\rm tot}[q,N_{c}]\right)\,,$ (42)
@@
$\bar{G}_{\rm NBC}(q_{f},p_{i})$ $=$ $\exp\biggl{[}\frac{i}{\hbar}\left\{p_{i}q_{f}-\frac{p_{i}^{2}}{2m}-\frac{\lambda\left(\lambda-3p_{i}+6mq_{f}\right)}{6m}\right\}\biggr{]}\,,$ (45)
@@
where $p_{i}$ is the initial momentum, $q_{i}$ in initial position, $P_{i}=p_{i}+\beta\,q_{i}$ and $Q_{f}=q_{f}$ is the final position. The details of this derivation can be found in the section 2 of Ailiga:2023wzl .
@@ III.1 NBC at $t=0$
It should be emphasized that the path-integral can also be performed using the zeta-function Narain:2022msz . However, the results differ from above by a numerical pre-factor. We believe that doing path-integral via time-slicing method is not only more rigorous and reliable but also fixes the normalization correctly.
@@
$G_{\rm NBC}[{\rm Bd}_{f},{\rm Bd}_{i}]=\frac{1}{2}\int_{-\infty}^{\infty}{\rm d}N_{c}\,\,\exp\left(\frac{i}{\hbar}S_{\rm tot}^{\rm on-shell}[\bar{q},N_{c}]\right)\,,$ (50)
where $S^{\rm on-shell}_{\rm tot}[\bar{q},N_{c}]$ is given in eq. (34). This can be done by doing a change of variables achieved by shifting the lapse $N_{c}$ by a constant
@@
$\Psi_{1}(\pi_{i})=\frac{1}{2}\exp\left[\frac{i}{\hbar}\left(\frac{3}{\Lambda}+4\alpha\right)\left(k\pi_{i}+\frac{\pi_{i}^{3}}{27}\right)\right]\,,$ (53)
@@
$G_{\rm NBC}[{\rm Bd}_{i},{\rm Bd}_{f}]=$ $\frac{\sqrt{3}}{2}\left(\frac{3\hbar}{\Lambda^{2}}\right)^{\frac{1}{3}}\exp\left[\frac{i}{\hbar}\left(\frac{3}{\Lambda}+4\alpha\right)\left(k\pi_{i}+\frac{\pi_{i}^{3}}{27}\right)\right]$ (55)
@@ III.2 RBC at $t=0$
$\int_{0^{+}}^{\infty}{\rm d}N_{c}\exp\left(\frac{3kN_{c}i}{\hbar}\right)\bar{G}_{\rm NBC}[q_{f},t=1;\tilde{p},t=0]=\biggl{.}\Psi_{1}(\tilde{p})\biggr{|}_{\alpha=0}\times\Psi_{2}(q_{f})\,,$ (58)
@@
$G_{\rm RBC}[{\rm Bd}_{i},{\rm Bd}_{f}]=$ $\sqrt{\frac{3\pi\hbar}{2i\beta}}\exp\left[\frac{4i\alpha}{\hbar}\left(kP_{i}+\frac{P_{i}^{3}}{27}\right)\right]\exp\biggl{[}\frac{i\bigl{(}-18k\beta^{2}+2P_{i}^{2}\beta^{2}+6P_{i}\beta\Lambda+3\Lambda^{2}\bigr{)}}{4\hbar\beta^{3}}\biggr{]}$ (60)
@@
It is noticed that for the RBC case, the path-integral once again consists of two parts: one dependent only on initial parameters ($P_{i}$ and $\beta$) and another only on $q_{f}$.
It is worth making the comparison with the NBC results in the limit $\beta\to 0$, in which limit the RBC reduces to NBC. To see this, we use the following identity
@@ IV.1 Implications of having Gauss-Bonnet
On extremization of Euclidean action $A_{E}(y)$ we observe that for $\alpha>0$ the point $y=-3\sqrt{k}$ corresponds to a stable saddle point leading to a positive exponent (aka Hartle-Hawking ) while $y=3\sqrt{k}$ corresponds to unstable saddle point leading to exponential with a negative argument (Vilenkin Tunneling). Coincidently, as we have shown in IV.2 and was also mentioned in DiTucci:2019bui ; Narain:2022msz , $P_{i}=-3i\sqrt{k}$ also corresponds to initial condition giving regular and well behaved geometries within a certain regime of $\beta$.
@@ IV.2.1 $0\leq x\leq 1$
This is actually the Hartle-Hawking state with positive weighting. The Gauss-Bonnet modification doesn’t disturb the structure of no-boundary weighting. The emergence of this structure further supports the findings of the paper Jonas:2020pos , where it was noticed that Higher-derivative corrections doesn’t destroy the no-boundary state of Universe and this is a room for the particular proposal to remain valid. The results obtained in Jonas:2020pos although perturbative still gives an assurance of the robustness of the no-boundary state of Universe. The topological nature of Gauss-Bonnet gravity provides extra freedom to go beyond perturbative study Ailiga:2023wzl . Here it is also seen that this exponential pre-factor is non only independent of the $\beta$-parameter but also hinting that lower value of $\Lambda$ is more favorable. This also covers the NBC case with $\beta=0$.
@@ IV.2.2 $x>1$
Then we notice that $f(1)=1$ and $f(\infty)=-1$. If we take derivative $f^{\prime}(x)=-6(x-1)^{2}/x^{4}0$, then it is seen that it is always negative for all values of $x$ except $x=1$ and $x=\infty$. It is a monotonically decreasing function of $x$ in the range $1$ and $-1$, but at some point $x_{0}$ the function becomes zero. This $x_{0}=2+2^{1/3}+2^{2/3}$ is a crossover point. The function $f(x)>0$ for $1<x\leq x_{0}$, and $f(x)<0$ for $x>x_{0}$.
@@ IV.3 Robin BC Interpretation
This is a Gaussian state with an imaginary momentum $P_{i}$ with the peak at $q_{i}=0$. It describes a state with shared uncertainty between the scale factor and the momentum of the universe at $t=0$. Taking this as an initial state DiTucci:2019bui ; DiTucci:2019dji , one can express the wave function as a gravitational path-integral in the following way
@@ V Picard-Lefschetz Methodology: mini review
The mini-superspace path-integral mentioned in eq. (42) involves an integration over the lapse $N_{c}$. Both in the NBC and RBC case one has to workout this integral either by directly computing it (as in NBC) or converting it into an integral over $\tilde{p}$ (as in RBC ). These integrals can also be worked out in saddle-point approximation and using Picard-Lefschetz methods. These methods helps us in analysing the behavior of the integrand in the complex plane Witten:2010cx ; Witten:2010zr ; Basar:2013eka ; Tanizaki:2014xba . This is achieved by determining the set of saddle points in the complex plane and finding the associated steepest ascent/descent paths. Then the Picard-Lefschetz methods systematically tell us the deformed integration contour along which the integrand is well-behaved and integral is convergent. To understand this in more detail we consider a generic oscillatory integral in complex plane
$I=\int{\cal D}z\,e^{i{\cal S}(z)/\hbar}\,,$ (77)
where the exponent is function of $z$. Here $\hbar$ is a real parameter, the action ${\cal S}(z)$ is a real-valued function. Generically, such integrands are highly oscillatory and hence the corresponding integration may not convergent along the given contour. However, it could be conditionally convergent in the sense that the integrand becomes well-behaved along the new contour of integration. Picard-Lefschetz (PL) methodology gives a systematic procedure for determining the deformed contour of integration. In PL theory one analytically continues both the integration parameter $z$ and the action ${\cal S}(z)$ into the complex plane. ${\cal S}(z)$ is then interpreted as an holomorphic functional of $z$. This implies that ${\cal S}$ satisfies a functional form of Cauchy-Riemann conditions

### v2

@@ II Mini-superspace action and Boundary Terms
Interestingly, for conformal-flat (Weyl-flat) metrics, there is an identity relating the Riemann tensor with the Ricci-tensor and the Ricci scalar. This is given by,
@@
This is a very useful relation when it is utilized in the GB-gravity action then, we get the following for case of Weyl-flat metrics
@@
where $V_{D-1}$ is the volume of $D-1$ dimensional sphere (for $\kappa=1$) and is given by  [^1] For $\kappa=(0,-1)$, the spatial slices are non-compact and $V_{D-1}$ wouldn’t be finite. In these cases, we have to either use an appropriate regularization scheme to regulate the volume or take non-trivial topologies.,
@@
Lapse and scale-factor rescaling allows one to recast this action into a simpler form. This change of variables is given by
@@
Under this transformation, the gravitational action becomes the following
@@
In this the $(\dot{})$ represent time $t$-derivative. For some terms, one can do the integration by parts which allows us to simplify a bit more and combine some of the terms. This eventually leads to
@@
Notice that there are two total derivative terms: one arises from the EH-part of gravitational action, while the other is the GB term itself which is topological in $4D$. We will keep $\Lambda$ to be arbitrary through the following sections till section IV, where we take it to be positive and do saddle point studies. Note also that in the following, we will make use of the convention $V_{3}=8\pi G_{N}$.
At this point, one can set up the variational problem by varying the action given in eq. (19) with respect to $q(t)$, which we do it by working in the ADM gauge $\dot{N}=0$ (setting $N(t)=N_{c}$, a constant). This can be implemented systematically by writing
@@
where one assumes $\bar{q}(t)$ to satisfy the equation of motion, $\delta q(t)$ is the fluctuation around this. The parameter $\epsilon$ is inserted to keep track of terms in expansion, which becomes handy as one goes to higher orders in perturbative expansion.
@@
Note that in the above there are two total time-derivative terms. These eventually become boundary terms. For a consistent variational problem, these terms must either vanish for the given choice of boundary conditions or need to be appropriately cancelled by the addition of extra boundary terms. On the other hand, the collection of terms proportional to $\delta q$ gives the equation of motion for $q$,
@@
$\delta S_{\rm bdy}=\frac{\epsilon}{2}\biggl{[}\frac{3}{N_{c}}\left(\bar{q}_{f}\delta\dot{q}_{f}-\bar{q}_{i}\delta\dot{q}_{i}\right)+24\alpha\left\{\left(\frac{k\delta\dot{q}_{f}}{2N_{c}}+\frac{\dot{\bar{q}}_{f}^{2}\delta\dot{q}_{f}}{8N_{c}^{3}}\right)-\left(\frac{k\delta\dot{q}_{i}}{2N_{c}}+\frac{\dot{\bar{q}}_{i}^{2}\delta\dot{q}_{i}}{8N_{c}^{3}}\right)\right\}\biggr{]}\,,$ (25)
@@
$\delta S_{\rm bdy}=-\epsilon\biggl{[}\left(\bar{q}_{f}\delta\pi_{f}-\bar{q}_{i}\delta\pi_{i}\right)+4\alpha\left\{\left(k\delta\pi_{f}+\frac{\bar{\pi}_{f}^{2}\delta\pi_{f}}{27}\right)-\left(k\delta\pi_{i}+\frac{\bar{\pi}_{i}^{2}\delta\pi_{i}}{27}\right)\right\}\biggr{]}\,.$ (28)
@@ II.2 Robin Boundary condition (RBC) at $t=0$
.$\biggl{.}\delta S_{\rm bdy}\biggr{|}_{\rm RBC}=-\epsilon\biggl{[}\bar{q}_{f}\delta\pi_{f}+\beta\bar{q}_{i}\delta q_{i}+4\alpha\left\{k\delta\pi_{f}+\frac{\bar{\pi}_{f}^{2}\delta\pi_{f}}{9}+\beta\left(k\delta q_{i}+\frac{(\bar{P}_{i}-\beta\bar{q}_{i})^{2}\delta q_{i}}{9}\right)\right\}\biggr{]}\,.$ (36)
@@ III Transition Amplitude
$G[{\rm Bd}_{f},{\rm Bd}_{i}]=\int_{\bf C}{\rm d}N_{c}\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}{\cal D}q(t)\,\,\exp\left(\frac{i}{\hbar}S_{\rm tot}[q,N_{c}]\right)\,,$ (42)
@@
$\bar{G}_{\rm NBC}(q_{f},t=1;p_{i},t=0)$ $=$ $\exp\biggl{[}\frac{i}{\hbar}\left\{p_{i}q_{f}-\frac{p_{i}^{2}}{2m}-\frac{\lambda\left(\lambda-3p_{i}+6mq_{f}\right)}{6m}\right\}\biggr{]}\,,$ (45)
@@
where $p_{i}$ is the initial momentum, $q_{i}$ in initial position, $P_{i}=p_{i}+\beta\,q_{i}$ and $Q_{f}=q_{f}$ is the final position. The details of this derivation can be found in section 2 of Ailiga:2023wzl .
@@ III.1 NBC at $t=0$
It should be emphasized that the path-integral can also be performed using the zeta-function Narain:2022msz . However, the results differ from above by a numerical pre-factor. It is worth noting that the quantum fluctuation determinant for the Neumann boundary condition results in a constant, which is 1, that agrees with the Van Vleck-Pauli-Morette determinant (see doi:10.1142/7305 ). For a more detailed analysis, see Appendix A.
@@
$G_{\rm NBC}[{\rm Bd}_{f},{\rm Bd}_{i}]=\int_{-\infty}^{\infty}{\rm d}N_{c}\,\,\exp\left(\frac{i}{\hbar}S_{\rm tot}^{\rm on-shell}[\bar{q},N_{c}]\right)\,,$ (50)
where $S^{\rm on-shell}_{\rm tot}[\bar{q},N_{c}]$ is given in eq. (34). By extending the range of integration all the way to $\infty$- ∞, one constructs the “ wave function ”, which is the solution of the WDW equation, as we later show in sec.VII. This is in contrast to the equation in 42, which computes the “propagator”. The integral can be done exactly by doing a change of variables achieved by shifting the lapse $N_{c}$ by a constant
@@
$\Psi_{1}(\pi_{i})=\exp\left[\frac{i}{\hbar}\left(\frac{3}{\Lambda}+4\alpha\right)\left(k\pi_{i}+\frac{\pi_{i}^{3}}{27}\right)\right]\,,$ (53)
@@
$G_{\rm NBC}[{\rm Bd}_{f},{\rm Bd}_{i}]=$ $\sqrt{3}\left(\frac{3\hbar}{\Lambda^{2}}\right)^{\frac{1}{3}}\exp\left[\frac{i}{\hbar}\left(\frac{3}{\Lambda}+4\alpha\right)\left(k\pi_{i}+\frac{\pi_{i}^{3}}{27}\right)\right]$ (55)
@@ III.2 RBC at $t=0$
$\int_{-\infty}^{\infty}{\rm d}N_{c}\exp\left(\frac{3kN_{c}i}{\hbar}\right)\bar{G}_{\rm NBC}[q_{f},t=1;\tilde{p},t=0]=\biggl{.}\Psi_{1}(\tilde{p})\biggr{|}_{\alpha=0}\times\Psi_{2}(q_{f})\,,$ (58)
@@
$G_{\rm RBC}[{\rm Bd}_{f},{\rm Bd}_{i}]=$ $\sqrt{\frac{6\pi\hbar}{i\beta}}\exp\left[\frac{4i\alpha}{\hbar}\left(kP_{i}+\frac{P_{i}^{3}}{27}\right)\right]\exp\biggl{[}\frac{i\bigl{(}-18k\beta^{2}+2P_{i}^{2}\beta^{2}+6P_{i}\beta\Lambda+3\Lambda^{2}\bigr{)}}{4\hbar\beta^{3}}\biggr{]}$ (60)
@@
It is worth noting that for the RBC case also, the path-integral once again factorizes into two parts: one depends only on initial parameters ($P_{i}$ and $\beta$), and another only on final $q_{f}$. This factorization was known previously to hold at the semiclassical level. However, at the exact expression, it was first shown in Ailiga:2023wzl .
For $\alpha>0$, a constraint motivated from Cheung:2016wjt ; Zwiebach:1985uq ; Gross:1986mw ; Metsaev:1987zx ; Chakravarti:2022zeq , it is noticed that the transition amplitude peaks around the initial parameter $P=-3i$, which is also the case, in absence of the Gauss-Bonnet for $\Lambda>0$. This result is illustrated in Figure 1, where we plot the term dependent on initial parameters, $G_{\rm RBC}(q_{f},P_{i})$, as a function of the Euclideanized variable $P_{i}=iy$. This favours the Hartle-Hawking no boundary geometry as the most dominant geometry in the path integral.
Figure 1: The plot of initial parameter dependent part of transition amplitude $G(q_{f},P_{i})$, given in Eq. (60). Here, we choose parameter values: $k=1$, $\Lambda=3$, $q_{f}=10$, $\hbar=0.0001$, $\alpha=1$ and $\beta=-1.5i$. The $G(q_{f},P_{i})$ axis has been scaled appropriately for ease of visualization.
In the subsequent section, we have shown it analytically. It is worth making the comparison with the NBC results in the limit $\beta\to 0$, in which limit the RBC reduces to NBC. To see this, we use the following identity
@@ IV.1 Implications of having Gauss-Bonnet
On extremization of Euclidean action $A_{E}(y)$ we observe that for $\alpha>0$ the point $y=-3\sqrt{k}$ corresponds to a stable saddle point leading to a positive exponent (aka Hartle-Hawking ) while $y=3\sqrt{k}$ corresponds to unstable saddle point leading to exponential with a negative argument (Vilenkin Tunneling). Coincidently, as we have shown in IV.2 and was also mentioned in DiTucci:2019bui ; Narain:2022msz , $P_{i}=-3i\sqrt{k}$ also corresponds to initial condition giving regular and well-behaved geometries within a certain regime of $\beta$. Since a universe with a higher positive value of the exponent is more probable, the choice $P_{i}=-3i\sqrt{k}$ is favored, as it maximizes the Gauss-Bonnet exponent.
Although, technically allowing $\alpha<0$ would result in the point $P_{i}=3i\sqrt{k}$ becoming a dominant geometry in the path integral, which would lead to a Vilenkin Tunneling proposal. However, the studies in string theory, such as those in Cheung:2016wjt ; Zwiebach:1985uq ; Gross:1986mw ; Metsaev:1987zx , as well as observational constraints from gravitational wave data and the area theorem, discussed in Chakravarti:2022zeq ; Isi:2020tac , limit the value of $\alpha$ to positive values, restrict this possibility.
@@ IV.2.1 $0\leq x\leq 1$
This is the Hartle-Hawking state with positive weighting ($\Lambda>0$). The Gauss-Bonnet modification doesn’t change the overall structure of the exponent. This further supports the findings of the paper Jonas:2020pos , where it was noticed that Higher-derivative corrections don’t prohibit the no-boundary Universe as a solution. The analysis done in Jonas:2020pos was perturbative in nature. However, the topological nature of Gauss-Bonnet gravity helps us to go beyond the perturbative studies Ailiga:2023wzl . Although the contribution from Gauss-Bonnet appears as a multiplicative factor to the Einstein-Hilbert gravity, it is necessary to correctly reproduce the gravitation entropy of HDG theory Shu:2008yd . It is known that the H-H exponent is equal to entropy/2 in Einstein’s gravity. Our study also establishes this to be true for the higher-derivative gauss-bonnet gravity. Here, it is also seen that this exponential pre-factor is not only independent of the $\beta$-parameter but also hint that a lower value of $\bar{\Lambda}$ is more favourable.
@@ IV.2.2 $x>1$
Then we notice that $f(1)=1$ and $f(\infty)=-1$. If we take derivative $f^{\prime}(x)=-6(x-1)^{2}/x^{4}<0$, then it is seen that it is always negative for all values of $x$ except $x=1$ and $x=\infty$. It is a monotonically decreasing function of $x$ in the range $1$ and $-1$, but at some point $x_{0}$ the function becomes zero. This $x_{0}=2+2^{1/3}+2^{2/3}$ is a crossover point. The function $f(x)>0$ for $1<x\leq x_{0}$, and $f(x)<0$ for $x>x_{0}$.
@@ IV.3 Robin BC Interpretation
This is a coherent state with an imaginary momentum $P_{i}$ with the peak at $q_{i}=0$. One must be a little bit careful before making such an interpretation. Substituting naively $P_{i}=-ip$ with $p$ being real in eq. 75, one ends up with a real state with zero (mean) momentum with a peak at $q_{i}=p/|\beta|$. However, we should view the state in eq. 75 as a complex generalization of the usual coherent state in the (analytically continued) phase space ($P_{i},q_{i}$). With this in mind, eq.75 describes a state with shared uncertainty between the scale factor and the momentum of the universe at $t=0$. Taking this as an initial state DiTucci:2019bui ; DiTucci:2019dji , one can express the wave function as a gravitational path-integral in the following way
@@ V Picard-Lefschetz Methodology: mini review
The mini-superspace path-integral mentioned in eq. (42) involves an integration over the real lapse $N_{c}$. Both in the NBC and RBC case one has to workout this integral either by directly computing it (as in NBC) or converting it into an integral over $\tilde{p}$ (as in RBC ). These integrals can also be worked out in saddle-point approximation and using Picard-Lefschetz methods. These methods helps us in analysing the behavior of the integrand in the complex plane Witten:2010cx ; Witten:2010zr ; Basar:2013eka ; Tanizaki:2014xba . This is achieved by determining the set of saddle points in the complex plane and finding the associated steepest ascent/descent paths. Then the Picard-Lefschetz methods systematically tell us the deformed integration contour along which the integrand is well-behaved and integral is convergent. To understand this in more detail, we consider a generic oscillatory integral
[…]
where the exponent is a function of $z$ (real). Here $\hbar$ is a real parameter, and the action ${\cal S}(z)$ is a real-valued function. Generically, such integrals are not absolutely convergent along the given contour. It may be either divergent or conditionally convergent. However, if it is conditionally convergent the Picard-Lefschetz (PL) methodology gives a systematic procedure for determining the deformed contour of integration along which the integral becomes absolutely convergent. In PL theory, one analytically continues both the integration parameter $z$ and the function ${\cal S}(z)$ into the complex plane. ${\cal S}(z)$ is then interpreted as an holomorphic function of $z$. This implies that ${\cal S}$ satisfies the Cauchy-Riemann conditions